### v1

@@ 5.2 Implementation details
For the $k$-SPC deep denoiser $\operatorname{D}_{\sigma}$ in $\operatorname{C}=\alpha(\operatorname{Id}-\operatorname{D}_{\sigma})$, we adopt the DRUNet architecture [54], following [50], with pretrained weights for grayscale and color images ($k=0.9$) from the GitHub repository [^7] https://github.com/FizzzFizzz/Learning-Pseudo-Contractive-Denoisers-for-Inverse-Problems. To simplify tuning and accelerate convergence, we fix the stepsize $\tau=1$, initial penalty parameter $\rho_{0}=10^{-4}$, speed factor $\nu=1.02$, and schedule the relaxation parameter $\lambda_{t}=1$ for $t<100$ and $\lambda_{t}=\frac{100}{t}$ for $t\geq 100$, satisfying the conditions $\sum\lambda_{t}=\infty$ and $\sum\lambda_{t}^{2}<\infty$ in Corollary 4.4. The source code is publicly available at the GitHub repository [^8] https://github.com/peterchen96/TensorCompletionMIP.
@@
with $\varphi>0$ and $\omega>1$. Since $f_{\varphi,\omega}$ is $\frac{1}{\omega-1}$-weakly convex, the GTCTV prior $\|\cdot\|_{\text{GTCTV}}$ is $\frac{4}{\omega-1}$-weakly convex on $\mathbb{V}$, by Lemma 3.3. The traffic data are reshaped into a tensor of dimensions (locations/sensors $\times$ time intervals × $\times\,1\,\times$ days), with $N_{\text{in}}=5$, $\Gamma=\{1,2,4\}$ and other hyperparameters set as follows:
@@
Table 2: Average quantitative results of different methods for MDI completion, with the best and second-best results highlighted in bold and underlined, respectively.
@@
Ours 27.511 0.809 29.698 0.873 32.817 0.931
5.3 Completion performence
In Table 2, we present the average quantitative results for MDI completion across various methods. Table 2 shows that Abs-TCTV-DPC performs particularly well on more intricate data, such as MSIs and color videos. Notably, at a low sampling rate of 0.05, Abs-TCTV-DPC achieves an average MPSNR improvement of 0.812 dB for MSIs and 0.654 dB for color videos over the second-best method, highlighting its superior completion performance.
In Fig. 3, we present the visual results for the color video bus at SR = 0.05. The visual results indicate that Abs-TCTV-DPC excels in preserving both the overall structural coherence and intricate local patterns, delivering clearer and more detailed reconstructions even at low sampling rates. Additionally, Fig. 4 illustrates the MPSNR curves for the MSI paints and the video bus across all methods. The curves illustrate that GTCTV-DPC achieves stable and progressive MPSNR improvements over iterations, highlighting its robustness and consistency.
Figure 3: Completion results for the color video bus (SR = 0.05), showing the 25th frame of each completed video. The MPSNR is indicated in the upper-left corner of each image.
Figure 4: (a) and (b): MPSNR curves with the $x$-axis denoting the iteration number for the MSI paint (a) and the video bus (b) at SR = 0.05. Note that for LRTFR and DRO-TFF, which run up to 3000 iterations, we record the MPSNR every 15 iterations. (c): Convergence performance within the monotone inclusion paradigm for the video bus at SR = 0.05, 0.10, and 0.20.
@@ 5.3 Completion performence
where $\operatorname{A}$, $\operatorname{B}$, and $\operatorname{C}$ correspond to the operators in Eq. Eq. 11. Here $\operatorname{A}=\partial\delta_{\mathcal{Y},\Omega}$ is the subdifferential of the data-consistency indicator; by Example 3.5 in [4] we have
@@
$=$ $\left\{\mathcal{X}\in\mathbb{V}\mid\langle\mathcal{X},\mathcal{Z}-\mathcal{X}_{t}^{\operatorname{A}}\rangle\leq 0,\forall\mathcal{Z}\in\mathbb{V}\text{ such that }\mathcal{P}_{\Omega}(\mathcal{Z})=\mathcal{P}_{\Omega}(\mathcal{Y})\right\}.$
We select $\mathcal{O}\in\operatorname{A}\mathcal{X}_{t}^{\operatorname{A}}$ as the subgradient satisfying this inclusion. For the GTCTV term, $\operatorname{B}\mathcal{X}_{t}^{\operatorname{A}}$ is obtained by evaluating a representative subgradient of the GTCTV penalty at $\mathcal{X}_{t}^{\operatorname{A}}$. In practice, we compute $\left\|\mathcal{X}_{t}^{\operatorname{A}}\right\|_{\text{GTCTV}}$ and employ the ‘torch.autograd‘ module to obtain its gradient automatically. Finally, $\operatorname{C}(\mathcal{X})=\alpha\big(\operatorname{Id}-\operatorname{D}_{\sigma}\big)(\mathcal{X})$ is evaluated directly using the DPC denoiser. Figure 4(c) demonstrates that $\operatorname{Tol}_{\mathrm{MIP}}(t)$ decays to near zero as iterations progress, indicating convergence of the iterates to a solution of the monotone inclusion problem in Eq. Eq. 12. Notably, these experiments incorporate the practical acceleration heuristics outlined in Algorithm 2 and Section 5.2. Despite these speedups, the monotone inclusion residual converges, providing empirical support for the theoretical global convergence result.
Table 3: The quantitative results of different methods for traffic data completion.
@@
Table 3 summarizes the quantitative results across multiple methods. The results indicate that SCAD-TCTV-DPC achieves competitive imputation performance compared to both low-rank methods with nonconvex penalties and multi-prior approaches, particularly at a low sampling rate of 0.30. Figure 5 presents the imputation results for two selected days from PeMS at SR = 0.30, demonstrating the superior performance of our method in effectively preserving the global structure while maintaining intricate local details of the original traffic data.
In section SM2 of the supplement, we further evaluate our method on color image completion in subsection SM2.2, present ablation studies in subsection SM2.3, describe the hyper-parameter tuning strategy in subsection SM2.4, and analyze computational performance in subsection SM2.5. Additional visual results are provided in subsection SM2.6.
6 Conclusions
@@
In this work, we present a novel tensor completion method within the monotone inclusion paradigm. To effectively capture global structure, we generalize the t-CTV prior with a weakly convex penalty and rigorously established its weak convexity in Lemma 3.3. To preserve intricate local details, we incorporate DPC denoisers [50] and establish their connection with $\beta$-cocoercivity in Lemma 3.4. Leveraging the DYS scheme, we derive the GTCTV-DPC method in Algorithm 2.
@@
A key theoretical contribution is our analysis of DYS for the general MIP in Eq. Eq. 7. Unlike previous studies [1, 16], we demonstrate that the associated operator $\operatorname{T}$ in Eq. Eq. 8 is SPC and extend the admissible stepsize range, together with distinct conditions for the relaxation parameters. This yields Corollary 4.4, which guarantees global convergence of Algorithm 2 to a solution of the proposed model Eq. Eq. 12. Empirical results on MDI and traffic datasets further demonstrate the superior performance and strong visual fidelity of the proposed method.
@@
Despite these advances, the reliance on multiple t-SVD operations results in high computational complexity, particularly for large-scale tensors. Future research will investigate randomized SVD algorithms [19] to enhance computational efficiency and investigate the adaptability of our framework to a wider range of tensor recovery tasks with tailored weakly convex penalties to further enhance practical applicability.

### v2

@@ IV-B Implementation details
For the $k$-SPC deep denoiser $\operatorname{D}_{\sigma}$ in $\operatorname{C}=\alpha(\operatorname{Id}-\operatorname{D}_{\sigma})$, we adopt the DRUNet architecture [51], following [9], with pretrained weights for grayscale and color images ($k=0.9$) from the GitHub repository [^7] https://github.com/FizzzFizzz/Learning-Pseudo-Contractive-Denoisers-for-Inverse-Problems. To simplify tuning and accelerate convergence, we fix the stepsize $\tau=1$, initial penalty parameter $\rho_{0}=10^{-4}$, speed factor $\nu=1.02$, and schedule the relaxation parameter $\lambda_{t}=1$ for $t<100$ and $\lambda_{t}=\frac{100}{t}$ for $t\geq 100$, satisfying the conditions $\sum\lambda_{t}=\infty$ and $\sum\lambda_{t}^{2}<\infty$ in Corollary 8.
@@
with $\varphi>0$ and $\omega>1$. Since $f_{\varphi,\omega}$ is $\frac{1}{\omega-1}$-weakly convex, the GTCTV prior $\|\cdot\|_{\text{GTCTV}}$ is $\frac{4}{\omega-1}$-weakly convex on $\mathbb{V}$, by Lemma 3. The traffic data are reshaped into a tensor of dimensions (locations/sensors $\times$ time intervals × $\times\,1\,\times$ days), with $N_{\text{in}}=5$, $\Gamma=\{1,2,4\}$ and other hyperparameters set as follows:
@@
TABLE II: Average quantitative results of different methods for MDI completion, with the best and second-best results highlighted in bold and underlined, respectively.
@@
Ours 27.506 0.809 29.694 0.873 32.813 0.931
IV-C Completion performence
In Table II, we present the average quantitative results for MDI completion across various methods. Table II shows that Abs-TCTV-DPC performs particularly well on more intricate data, such as MSIs and color videos. Notably, at a low sampling rate of 0.05, Abs-TCTV-DPC achieves an average MPSNR improvement of 0.717 dB for MSIs and 0.649 dB for color videos over the second-best method, highlighting its superior completion performance.
In Fig. 3, we present the visual results for the color video stefan at SR = 0.05. The visual results indicate that Abs-TCTV-DPC excels in preserving both the overall structural coherence and intricate local patterns, delivering clearer and more detailed reconstructions even at low sampling rates. Additionally, Fig. 4 illustrates the MPSNR curves for the MSI paints and the video bus at SR = 0.05 across all competing methods. The curves illustrate that GTCTV-DPC achieves stable and progressive MPSNR improvements over iterations, highlighting its robustness and consistency.
Figure 4: (a) and (b): MPSNR curves with the $x$-axis denoting the iteration number for the MSI paints (a) and the video bus (b) at SR = 0.05. Note that for LRTFR and DRO-TFF, which run up to 3000 iterations, we record the MPSNR every 15 iterations. (c): Convergence performance within the monotone inclusion paradigm for the video bus at SR = 0.05, 0.10, and 0.20.
@@ IV-C Completion performence
where $\operatorname{A}$, $\operatorname{B}$, and $\operatorname{C}$ correspond to the operators in Eq. (9), using the color video bus at SR = 0.05, 0.10, and 0.20 as examples. Here $\operatorname{A}=\partial\delta_{\mathcal{Y},\Omega}$ is the subdifferential of the data-consistency indicator; by Example 3.5 in [55] we have
@@
$=$ $\left\{\mathcal{X}\in\mathbb{V}\mid\langle\mathcal{X},\mathcal{Z}-\mathcal{X}_{t}^{\operatorname{A}}\rangle\leq 0,\right$.
$\left.\forall\mathcal{Z}\in\mathbb{V}\text{, such that }\mathcal{P}_{\Omega}(\mathcal{Z})=\mathcal{P}_{\Omega}(\mathcal{Y})\right\}.$
We select $\mathcal{O}\in\operatorname{A}\mathcal{X}_{t}^{\operatorname{A}}$ as the subgradient satisfying this inclusion. For the GTCTV term, $\operatorname{B}\mathcal{X}_{t}^{\operatorname{A}}$ is obtained by evaluating a representative subgradient of the GTCTV penalty at $\mathcal{X}_{t}^{\operatorname{A}}$. In practice, we compute $\left\|\mathcal{X}_{t}^{\operatorname{A}}\right\|_{\text{GTCTV}}$ and employ the ‘torch.autograd‘ module to obtain its gradient automatically. Finally, $\operatorname{C}(\mathcal{X})=\alpha\big(\operatorname{Id}-\operatorname{D}_{\sigma}\big)(\mathcal{X})$ is evaluated directly using the DPC denoiser. Fig. 4(c) demonstrates that $\operatorname{Tol}_{\mathrm{MIP}}(t)$ decays to near zero as iterations progress, indicating convergence of the iterates to a solution of the monotone inclusion problem in Eq. (10). Notably, these experiments incorporate the practical acceleration heuristics outlined in Algorithm 2 and section IV-B. Despite these speedups, the monotone inclusion residual converges, providing empirical support for the theoretical global convergence result.
TABLE III: The quantitative results of different methods for traffic data completion.
@@
Table III summarizes the quantitative results across multiple methods. The results indicate that SCAD-TCTV-DPC achieves competitive imputation performance compared to both low-rank methods with nonconvex penalties and multi-prior approaches, particularly at a low sampling rate of 0.30. Fig. 5 presents the imputation results for two selected days from PeMS at SR = 0.30, demonstrating the superior performance of our method in effectively preserving the global structure while maintaining intricate local details of the original traffic data. In section S.II.B of the supplement, we further evaluate our method on color image completion. Additional visual results are provided in section S.II.C.
IV-D Discussions
IV-D1 Ablation studies
We perform two ablation studies: (i) to evaluate the contributions of the GTCTV prior and the DPC denoiser in GTCTV-DPC for MDI completion; and (ii) to assess the effect of the SCAD penalty for spatio-temporal traffic-data completion. For the MDI completion ablation, we consider three variants of GTCTV-DPC:
  • TNN-DPC: Replace the GTCTV prior with TNN to test the effectiveness of GTCTV, while retaining the DPC denoiser;
  • Abs-TCTV-DNE: GTCTV with $f(x)=|x|$ paired with a deep non-expansive (Definition 5, NE) denoiser to test the effect of the denoiser assumption;
  • Abs-TCTV-DFNE: GTCTV with $f(x)=|x|$ and a deep firm non-expansive (Definition 4, FNE) denoiser [46].
These variants are tested on five YUV color videos (bus, mobile, akiyo, mother-daughter, tempete) at sampling rates of 0.05, 0.10, and 0.20. Hyperparameters follow GTCTV-DPC for color videos ($N_{\text{in}}=5$, $\Gamma=\{1,2,4\}$, $\sigma_{0}=0.30$, $\alpha=0.50$), except for TNN-DPC, which uses an initial TNN threshold of $1$ and divides the speed factor $\nu=1.02$ by the iteration number. All denoisers (DPC, NE, FNE) use the DRUNet architecture [51], with pretrained weights for NE [^8] https://github.com/FizzzFizzz/New-baseline-for-DRUNet-under-different-assumptions and FNE [^9] https://github.com/basp-group/PnP-MMO-imaging. Completion performance is evaluated using MPSNR and MSSIM, with results averaged over five runs reported in Table IV.
TABLE IV: Average MPSNR and MSSIM for GTCTV-DPC variants on five YUV color videos at sampling rates 0.05, 0.10, and 0.20.
Sampling Rate 0.05 0.10 0.20
Method MPSNR MSSIM MPSNR MSSIM MPSNR MSSIM
Observation 7.503 0.027 7.738 0.043 8.249 0.078
TNN-DPC 23.480 0.633 26.985 0.774 30.651 0.883
Abs-TCTV-DNE 26.869 0.781 29.365 0.856 32.780 0.924
Abs-TCTV-DFNE 26.442 0.765 28.943 0.843 32.341 0.917
Abs-TCTV-DPC 27.274 0.796 29.769 0.868 33.058 0.931
Figure 6: The results of color video completion on bus (SR = 0.05) by different variants of the proposed method.
Table IV shows that Abs-TCTV-DPC outperforms other variants across all sampling rates, with Abs-TCTV-DNE achieving the second-best performance. Fig. 6 shows the visual comparisons for the bus video (SR = 0.05). The GTCTV prior, incorporating the spatial and temporal gradient information, captures the holistic tensor structure more effectively than TNN, leading to higher MPSNR and MSSIM. Furthermore, as discussed in section II-B, the hierarchy of denoiser assumptions (FNE $\implies$ NE $\implies$ PC, Eq. (3)) indicates that the less restrictive PC assumption in GTCTV-DPC enables better completion performance compared to the stronger NE and FNE assumptions, as evidenced by the results.
TABLE V: The quantitative results for traffic data completion.
Sampling Rate 0.30 0.50 0.70
Method MAPE RMSE MAPE RMSE MAPE RMSE
Guangzhou $(214\times 144\times 61)$
Abs-TCTV-DPC 9.25 4.02 7.29 3.20 6.38 2.82
SCAD-TCTV-DPC 7.95 3.47 6.98 3.07 6.38 2.81
Seattle $(323\times 288\times 28)$
Abs-TCTV-DPC 7.43 4.68 6.59 4.27 6.15 4.03
SCAD-TCTV-DPC 5.95 3.73 5.28 3.37 4.87 3.14
PeMS $(228\times 288\times 44)$
Abs-TCTV-DPC 5.02 3.65 3.58 2.62 2.92 2.10
SCAD-TCTV-DPC 4.29 3.13 3.38 2.44 2.88 2.05
To assess the significance of the SCAD penalty in GTCTV-DPC for spatio-temporal traffic data completion, we compare SCAD-TCTV-DPC against Abs-TCTV-DPC. We test these variants on the Guangzhou ($214\times 144\times 61$), Seattle ($323\times 288\times 28$), and PeMS ($228\times 288\times 44$) datasets at sampling rates of 0.30, 0.50, and 0.70. Completion performance is evaluated using MAPE and RMSE, with results reported in Table V. Table V shows that SCAD-TCTV-DPC outperforms Abs-TCTV-DPC across all datasets and sampling rates, with lower MAPE and RMSE. The nonconvex SCAD penalty promotes sparsity and captures complex spatio-temporal patterns in traffic data more effectively than the convex absolute value function, demonstrating its suitability for such datasets.
Figure 7: Boxplot of average MPSNR for each $\sigma_{0}$-$\alpha$ combination for Abs-TCTV-DPC on the YUV color video subset (bus, mobile, akiyo) across sampling rates 0.05, 0.10, and 0.20. The selected combination ($\sigma_{0}=0.30$, $\alpha=0.50$) is highlighted.
IV-D2 Hyper-parameters tuning strategy
As introduced in section IV-A, for each dataset type with a fixed sampling rate, we choose the same small samples from selected datas, and utilize the grid search to find the optimal hyper-parameters for the small samples, and finally adopt the optimal hyper-parameters for all selected datas for all compared methods. Here we take the completion of color videos as an instance to introduce the procession of hyper-parameters tunning strategy for GTCTV-DPC.
For color videos, we select bus, mobile, and akiyo from the YUV dataset, as shown in Fig. S.2 from the supplement, and perform grid search at sampling rates of 0.05, 0.10, and 0.20. For GTCTV-DPC, two hyper-parameters need to be carefully tuned: the initial denoising strength $\sigma_{0}$ and the weighting factor $\alpha$ in the DPC denoiser $\operatorname{C}=\alpha(\operatorname{Id}-\operatorname{D}_{\sigma})$. We search over $\sigma_{0}\in\{0.20,0.30,0.50,0.70\}$ and $\alpha\in\{0.30,0.50,0.70\}$, resulting in 12 combinations. Fig. 7 presents a boxplot of the average MPSNR for each combination of $\sigma_{0}$ and $\alpha$ across all sampling rates, illustrating the robustness of GTCTV-DPC’s completion performance to different hyper-parameter choices. For color videos, we select $\sigma_{0}=0.30$ and $\alpha=0.50$, as these values balance high MPSNR and MSSIM with computational efficiency. These selected hyper-parameters were then applied to all color video datasets in our experiments.
TABLE VI: Computational time (mean $\pm$ std, in seconds).
Method Used Time (s)
HIR-Diff 63.25 $\pm$ 8.42
LRTFR 20.84 $\pm$ 0.31
DRO-TFF 21.65 $\pm$ 0.47
FTNN 10994.82 $\pm$ 107.80
t-CTV 1540.58 $\pm$ 5.85
DP3LRTC 1207.08 $\pm$ 11.13
FBGND 5543.76 $\pm$ 9.29
Ours 2125.01 $\pm$ 4.05
IV-D3 Computational performance
To test computational performance, we use the akiyo color video ($288\times 352\times 3\times 50$) from the YUV dataset with a sampling rate of 0.20. Each method is run independently five times, and Table VI reports the mean and standard deviation of computational time (in seconds). Fig. 8 shows the residual error (tolerance) as a function of iteration number for each method. For LRTFR and DRO-TFF, which are run for up to 3000 iterations, the tolerance is recorded every 15 iterations.
Figure 8: Residual error (tolerance) versus iteration number for compared methods on the akiyo video ($288\times 352\times 3\times 50$, sampling rate 0.20). For LRTFR and DRO-TFF, tolerance is recorded every 15 iterations due to their 3000-iteration limit.
Self-supervised deep-learning methods (HIR-Diff, LRTFR, DRO-TFF) leverage lightweight architectures and GPU acceleration, yielding high computational efficiency. However, their completion performance, as measured by MPSNR and MSSIM, remains inferior to that of GTCTV-DPC and t-CTV. The proposed GTCTV-DPC method, which integrates the GTCTV prior, exhibits rapid convergence in terms of residual error, as illustrated in Fig. 8. Nevertheless, due to the multiple t-SVD computations required per iteration, it incurs higher computational costs. Despite this, GTCTV-DPC achieves consistently superior completion performance, as shown in section IV-C. Moreover, at a lower sampling rate of 0.05, GTCTV-DPC demonstrates further improvement in completion quality, indicating its strong adaptability under highly undersampled conditions. In the future, we will explore the use of randomized SVD techniques to accelerate t-SVD computations and enhance scalability.
V Conclusions
@@
In this work, we present a novel tensor completion method within the monotone inclusion paradigm. To effectively capture global structure, we generalize the t-CTV prior with a weakly convex penalty and rigorously established its weak convexity in Lemma 3. To preserve intricate local details, we incorporate DPC denoisers [9] and establish their connection with $\beta$-cocoercivity in Lemma 4. Leveraging the DYS scheme, we derive the GTCTV-DPC method in Algorithm 2.
@@
A key theoretical contribution of this work lies in the convergence analysis of the GTCTV-DPC method. Starting from the general MIP in Eq. (6), we showed that the associated operator $\operatorname{T}$ in Eq. (12) is SPC and extended the admissible stepsize range, along with explicit conditions for relaxation parameters. This yields Corollary 8, which establishes the global convergence of Algorithm 2 to a solution of the proposed model in Eq. (10).
Empirical results on MDI and traffic datasets further demonstrate the superior performance and strong visual fidelity of the proposed method. For instance, at a sampling rate of 0.05 for MDI completion, GTCTV-DPC achieves an average MPSNR that surpasses the second-best method by 0.717 dB, and 0.649 dB for MSIs, and color videos, respectively.
@@
Despite these advances, the reliance on multiple t-SVD operations results in high computational complexity, particularly for large-scale tensors [73], as discussed in section IV-D3. Future research will investigate randomized SVD algorithms [74] to enhance computational efficiency and investigate the adaptability of our framework to a wider range of tensor recovery tasks with tailored weakly convex penalties to further enhance practical applicability.